OpenStudy (anonymous):

The shadow of a 5-foot tall post is 20 feet long. At the same time a tree casts a shadow that is 56 feet long. What is the height of the tree?

OpenStudy (anonymous):

\[20/5 = x/56\] what is your x?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

x =16

OpenStudy (anonymous):

got that backwards, \[5/20 = x/56\]

OpenStudy (anonymous):

x =14

OpenStudy (anonymous):

right, it becomes a simple ratio. the shadow of the man is 4 times longer than he is, thus the shadow of the tree is 4 times longer that it is.